OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (3, 4) and (–3, –8) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (helder_edwin):

first compute the slope do u know this?

OpenStudy (anonymous):

\[Slope = \frac{y_2 - y_1}{x_2 - x_1}\]

OpenStudy (anonymous):

Can you find slope from this??

OpenStudy (anonymous):

Here: \[(x_1, y_1) = (3,4)\] \[(x_2, y_2) = (-3, -8)\] Just plug in the values and find slope first..

OpenStudy (anonymous):

Yes is it m= 2

OpenStudy (anonymous):

Yes..

OpenStudy (waleed_imtiaz):

U can apply two-point form..... If U want to apply point slope form. then just calculate the slope first by the formula, it will be slope=2 and then use any of the point in the point slope form because these two points lie on the same line so u can put any point.........

OpenStudy (anonymous):

okay so the point slope is y - 4 = 2(x-3)

OpenStudy (anonymous):

Yes..

OpenStudy (waleed_imtiaz):

the equation would be 2x-y-2=0

OpenStudy (anonymous):

Okay now how do I explain part 2

OpenStudy (anonymous):

Because in both the cases we will have same y intercept.. It does not depend upon the points we have chosen..

OpenStudy (anonymous):

Okay thank you !

OpenStudy (waleed_imtiaz):

If u place (3,4), the equation would be 2x-y-2=0 and If U place (-3,-8), then the equation would be 2x-y+2=0. so u can see, only the constant term has the sign difference... and there's a supposition that if the co-efficients of x and y are same for both the equations and also have the same signs., then the equations are parallel, so U have the proof in these equations........